A car traveling at 50 metre per seconds from rest covers a distance of 10km calculate the acceleration
1
Expert's answer
2020-07-15T09:35:36-0400

Assuming the constant acceleration, the covered distance is given by:


"d = \dfrac{at^2}{2}"

where "d = 10\times10^3m" is the distance, "a" is the acceleration and "t" is the time.

On the other hand, by definition, the acceleration is:


"a = \dfrac{v - 0}{t} = \dfrac{v }{t}"

where "v = 50m\/s" is the final velocity and stands for the intitial velocity (start from rest).

Substituting the expression for "a" into the first equation and expressing "t" , get:


"d = \dfrac{vt^2}{2t} = \dfrac{vt}{2}\\
t = \dfrac{2d}{v}"

Substituting the last equation into the expression for "a", obtain:


"a = \dfrac{v^2}{2d} = \dfrac{50^2}{2\cdot 10\cdot 10^3} = 0.125 m/s^2"

Answer. 0.125 m/s^2.